Solve the following systems by substitution.
8x - 4y = 16
y = 2x - 4

Answer by ptaylor(2198)


OK. From Eq 2, y=2x-4. So, lets substitute 2x-4 for y in Eq 1:
8x-4(2x-4)=16 get rid of parens
8x-8x+16=16 subtract 16 from both sides:
0=0
NO SOLUTION EXISTS. WHY??????????? Eq 1 and Eq 2 are identical equations
First, we'll multiply Eq 2 by 4 and we get:
4y=8x-16 rearrange this equation by subtracting 4y and adding 16 to both sides
8x-4y=16-------------SAME AS EQ 1

Hope this helps---ptaylor